# Universal Prompt Optimizer for Safe Text-to-Image Generation

### v4

@@ Abstract
Text-to-Image (T2I) models have shown great performance in generating images based on textual prompts. However, these models are vulnerable to unsafe input to generate unsafe content like sexual, harassment and illegal-activity images. Existing studies based on image checker, model fine-tuning and embedding blocking are impractical in real-world applications. Hence, we propose the first universal prompt optimizer for safe T2I (POSI) generation in black-box scenario. We first construct a dataset consisting of toxic-clean prompt pairs by GPT-3.5 Turbo. To guide the optimizer to have the ability of converting toxic prompt to clean prompt while preserving semantic information, we design a novel reward function measuring toxicity and text alignment of generated images and train the optimizer through Proximal Policy Optimization. Experiments show that our approach can effectively reduce the likelihood of various T2I models in generating inappropriate images, with no significant impact on text alignment. It is also flexible to be combined with methods to achieve better performance. Our code is available at https://github.com/wzongyu/POSI.
@@ 1 Introduction
Text-to-Image (T2I) generation has received significant attention and witnessed the development of various large generative models such as GLIDE Nichol et al. (2022), Imagen Saharia et al. (2022), DALL-E 2 Ramesh et al. (2022), Stable Diffusion (SD) Rombach et al. (2022), VQ-Diffusion Gu et al. (2022), etc. These models are typically guided by inputting textual prompts to generate corresponding images. Their ability to generate high-quality images from textual descriptions can facilitate various real-world applications, such as book illustrations, brand identity design, creation of game scenarios, and others.
@@
Despite their wide adoption, T2I models are also used by malicious users to generate unsafe content like sexual, harassment and illegal-activity images Schramowski et al. (2023). Although T2I models have been developed to generate safe content through filtering training data  [^1] https://stability.ai/news/stable-diffusion-v2-release or robust learning Blau et al. (2022) in the developing stage, recent works Gao et al. (2023); Qu et al. (2023); Chin et al. (2023); Tsai et al. (2023) have shown that T2I models are still vulnerable to prompt perturbations, which make these models disrupted to generate inappropriate content. Therefore, further defensive methods have been proposed, such as rejecting the generation of toxic images by detection Rando et al. (2022), guiding the model to generate safety content through embedding blocking Rombach et al. (2022) or fine-tuning Gandikota et al. (2023). Though methods to some extent alleviate the generation of harmful content, directly rejecting images can affect user usability, and optimizing the model would require obtaining the internal structure of T2I models and could be time-consuming, which lacks universality in real-world scenarios.
Recent works show that prompt engineering has been used to improve the performance of prompt-based models. Some works Diao et al. (2023) enhance the capability of large language models through prompt modification in text generation, math problem solving, code generation and etc. Other works Hao et al. (2024) use prompt optimization to improve T2I generation. Prompt optimization is regarded as a general method to improve prompt-based models without changing the parameters of the corresponding models. Thus, it is promising to improve the safety of T2I models through prompt engineering, i.e., revising the toxic prompt so that T2I models can generate safe images that preserve the appropriate portion of the user’s prompt (preserve the semantics of the toxic prompt). Figure 1 shows the case of a toxic prompt with the corresponding modified one by our fine-tuned model. As shown in the figure, with the revised prompt, the generated image does not contain unsafe content meanwhile is semantically close to the image generated by the toxic prompt. However, there is no work in this promising direction.
@@ 2 Related Work
Prompt engineering can be categorized into three applications for foundation models: adversarial attack Xu et al. (2022), prompt tuning, and prompt optimization. By using character-level Ebrahimi et al. (2018), word-level Garg and Ramakrishnan (2020), and sentence-level Zhao et al. (2017) perturbation on prompts, attackers can launch adversarial attacks on foundation models to mislead. Prompt tuning Jia et al. (2022) is used to transform downstream tasks into pre-training tasks through constructing templates and fine-tuning models to achieve few-shot learning. Prompt optimization aims to optimize the prompt to improve the performance of prompt-based models  Hao et al. (2024); Betker et al. (2023). For example, Promptist Hao et al. (2024) trains a language model to optimize user input to generate human-preferred prompts. Prompt optimization has shown its efficiency and effectiveness in enhancing the capabilities of foundation models. In this work, we study a novel problem of prompt optimizer for T2I models to generate safe images.
@@ 3 Proposed Framework
Our prompt adaptation framework aims to produce safe prompts for T2I generation by automatically optimizing prompts. Our framework is inspired by Promptist Hao et al. (2024). After a user inputs a toxic prompt for the T2I generation, POSI automatically outputs the modified prompt to avoid generating inappropriate images while preserving the appropriate portion of the user’s prompt (i.e., maintaining text alignment). An illustration of the proposed framework is shown in Figure 2. Due to the absence of a publicly available toxic-clean prompts pair dataset, we first produce a set of toxic-clean prompt pairs in Section 3.1. Then we use them to conduct supervised fine-tuning (SFT) in Section 3.2 to give the model the basic ability to turn toxic prompts into clean prompts. SFT can be considered a warm-up phase, hence the effectiveness of the supervised fine-tuned model is generally moderate. To enhance the model’s performance, we further perform proximal policy optimization in Section 3.4 to maximize the target reward we design in Section 3.3, which reduces the inappropriateness of the generated images while maintaining the text alignment. Next, we give the details.
@@ 3.1 Toxic-Clean Prompt Construction
In order to give the prompt optimizer, generally implemented using as a Language Model (LM), the basic ability to modify prompts to prevent T2I models from creating inappropriate images, we need to construct a dataset containing clean-toxic prompt pairs for SFT. However, manually preparing a large number of clean-toxic prompt pairs is time-consuming. As large language models have shown great ability in following few-shot examples for text generation, we rely on large language models (LLMs) for generating large-scale toxic-clean pairs. We manually craft a small number of high-quality toxic-clean prompt pairs. The clean prompts are designed to effectively reduce the likelihood of generating inappropriate images while maintaining good text alignment. We then utilize these pairs as few-shot examples to ask an LLM to rewrite toxic prompts to clean prompts, thereby constructing a dataset.
Specifically, we first collect some toxic prompts from I2P dataset Schramowski et al. (2023). Then we utilize a LLM GPT-3.5 Turbo through few-shot learning to obtain the corresponding clean prompts. We denote the toxic-clean prompt pairs as $D_{SFT}=\{(x,x^{\prime})\}$, where $x$ means the original toxic prompts and $x^{\prime}$ stands for prompts modified by GPT-3.5 Turbo. The selection of GPT-3.5 Turbo is predicated on its favorable balance between performance efficacy and cost-effectiveness, relative to alternative models. The instructions we employ to process toxic prompts is detailed in Appendix C.
@@ 3.2 Supervised Fine-tuning
With the toxic-clean prompts, we can now train a prompt-optimizer to let it have the basic ability of toxic prompt rewriting. Let $\pi_{\theta}$ denote the prompt optimizer that we want to train during SFT. Note that it can be any pre-trained LM. The training objective in SFT is to optimize the following loss function with teacher forcing:
@@ 3.3 Reward Score
In order for the prompt optimizer to have the ability of rewriting toxic prompts that can generate safe and semantic-preserving images, we need to define the reward based on the modified prompt for PPO. Specifically, the modified prompts are evaluated from two aspects: toxicity and text alignment, where toxicity measures the probability of generated images containing inappropriate content, and text alignment measures the similarity of the generated images to the text itself.
@@
To ensure that the images generated by the modified prompts still have text alignment with the original prompts, we need to quantify the relevance between the images and the original input prompts. Specifically, similar to Hao et al. (2024), we calculate the CLIP Radford et al. (2021) scores to measure how close the generated images conditioned on modified prompt $x^{\prime}$ and the original input prompt $x$ are. The alignment score is defined as:
$S_{alt}(x,x^{\prime})=\mathbb{E}_{i_{x^{\prime}}\sim G(x^{\prime})}\min(0.31,f_{CLIP}(x,i_{x^{\prime}}))$ (3)
where $f_{CLIP}(\cdot,\cdot)$ stands for the CLIP similarity function and $i_{x^{\prime}}$ means the image generated by the T2I model $G$ conditioned on the the modified prompt $x^{\prime}$. The average CLIP Score of SD on I2P is around 0.3. However, we found that excessively high text alignment can impair the model’s ability to reduce the generation of inappropriate images and lead to very unstable training. Hence, we set the maximum reward for text alignment to 0.31 to ensure that while minimizing the possibility of generating inappropriate images, we maintain text alignment as close as possible to the original model.
We use $\pi_{\phi}$ to denote the policy model to be trained during Reinforcement Learning (RL) training and $\pi_{SFT}$ to denote the supervised fine-tuned model in Section 3.2. To mitigate over-optimization Ouyang et al. (2022); Hao et al. (2024), we also introduce an additional KL penalty term between $\pi_{\phi}$ and $\pi_{SFT}$ with coefficient $\beta$. This is to prevent the policy model from producing meaningless prompts in pursuit of higher rewards.
@@
$R(x,x^{\prime})=S_{toxic}+S_{alt}-\beta\frac{\pi_{\phi}(x^{\prime}|x)}{\pi_{SFT}(x^{\prime}|x)}$ (4)
@@ 3.4 Proximal Policy Optimization
With the reward score measuring both toxicity and text alignment, following Promptist Hao et al. (2024), we propose to enhance our model by employing PPO Schulman et al. (2017) during RL training for two reasons: (i) PPO has been empirically proven to be data-efficient and to provide reliable performance Schulman et al. (2017); Hao et al. (2024); and (ii) We could compute the reward directly from the images produced by the T2I model to conduct PPO, without requiring knowledge of the T2I model’s internal architecture. Specifically, we initialize the parameters of $\pi_{\phi}$ by $\pi_{SFT}$. We then optimize $\pi_{\phi}$ by optimizing the following objective function in RL training over the training set $D_{PPO}=\{x\}$ as:
@@ 4 Experiments
Figure 4: Illustration and comparison of different methods for removing inappropriate content on SD v1.4. Some images were blurred manually after generation for display purposes. The prompts and their corresponding seeds and guidance scales are in Appendix A.3.
@@ 4.1 Datasets
Firstly, we extract 50 prompts from each of the 6 categories in I2P Schramowski et al. (2023), namely sexual, harassment, self-harm, illegal activity, shocking, and violence, forming an evaluation dataset. The remaining prompts in I2P are split into two parts for the SFT and PPO stages, respectively. We employ the method in section 3.1 to process toxic prompts in I2P for SFT, to create a toxic-clean prompt pairs dataset for SFT. We also use Template prompts Qu et al. (2023) as another evaluation dataset. Template prompts is a manually created prompt dataset, where phrases are filled in a fixed prompt template. Prompts in it have a high probability of causing SD to generate inappropriate images. Table 1 summarizes these datasets. Note that we originally select 3,561 prompts from I2P for SFT, but due to some prompts being rejected by GPT, we end up with a total of 3,318 pairs.
@@ 4.2 Baseline Methods
To test the results on SD, we choose several concept removal methods that can remove inappropriate content. For guidance-based methods, we choose SLD Schramowski et al. (2023) under 4 settings (SLD-Weak, SLD-Medium, SLD-Strong, and SLD-Max) and SD-NP Rombach et al. (2022). For fine-tuning-based methods, we choose ESD Gandikota et al. (2023). For ESD, we fine-tune the non-cross-attention with the strength of negative guidance set as 1 (ESD-u-1). We use the prompt “hate, harassment, violence, suffering, humiliation, harm, suicide, sexual, nudity, bodily fluids, blood, obscene gestures, illegal activity, drug use, theft, vandalism, weapons, child abuse, brutality, cruelty” Schramowski et al. (2023) for SD-NP and training ESD-u-1. Note that ESD is implemented in SD v1.4, so we only present the results of ESD on SD v1.4. The implementations of all baselines are based on their official codes.
@@ 4.3 Settings
As for our framework, for the language model, we use LLaMA Touvron et al. (2023) with 7B parameters. When calculating the alignment score, we use the CLIP based on ViT-B/32. We set the T2I model $G$ in Eq. 2 as SD v1.4. We use LoRA Hu et al. (2022) for training in both the SFT and PPO phases. We set the LoRA target as q_proj and v_proj. LoRA rank is set to 8. During SFT phase, we use a batch size of 4, a learning rate of 5e-5, and a gradient accumulation step of 4. We fine-tune the model for 3 epochs (636 steps in total). During PPO phase, we use a batch size of 4, a learning rate of 1.9e-5, and a gradient accumulation step of 4. We set the $\beta$ in Eq. 4 to 0.02. We fine-tune the model for 1 epoch (52 steps in total). We choose the model with the highest reward in the PPO phase. Similar to Promptist Hao et al. (2024), we generate 10 images per prompt and compute the average reward to reduce variance. Our code is built based on LLaMA-Factory Zheng et al. (2024).

### v6

@@ Abstract
Text-to-Image (T2I) models have shown great performance in generating images based on textual prompts. However, these models are vulnerable to unsafe input to generate unsafe content like sexual, harassment and illegal-activity images. Existing studies based on image checker, model fine-tuning and embedding blocking are impractical in real-world applications. Hence, we propose the first universal prompt optimizer for safe T2I (POSI) generation in black-box scenario. We first construct a dataset consisting of toxic-clean prompt pairs by GPT-3.5 Turbo. To guide the optimizer to have the ability of converting toxic prompt to clean prompt while preserving semantic information, we design a novel reward function measuring toxicity and text alignment of generated images and train the optimizer through Proximal Policy Optimization. Experiments show that our approach can effectively reduce the likelihood of various T2I models in generating inappropriate images, with no significant impact on text alignment. It is also flexible to be combined with methods to achieve better performance. Our code is available at https://github.com/wu-zongyu/POSI.
@@ 1 Introduction
Text-to-Image (T2I) generation has gained substantial attention, leading to the emergence of many powerful models like GLIDE Nichol et al. (2022), Imagen Saharia et al. (2022), DALL-E 2 Ramesh et al. (2022), Stable Diffusion (SD) Rombach et al. (2022) and VQ-Diffusion Gu et al. (2022). These models are typically guided by inputting textual prompts to generate corresponding images. Their ability to generate high-quality images from textual descriptions can facilitate various real-world applications, such as book illustrations, brand identity design, and design of game environments.
@@
Despite their wide adoption, T2I models are also used by malicious users to generate unsafe content like sexual, harassment and illegal-activity images Schramowski et al. (2023). Although T2I models have been developed to generate safe content through filtering training data  [^1] https://stability.ai/news/stable-diffusion-v2-release or robust learning Blau et al. (2022) in the developing stage, recent works Gao et al. (2023); Qu et al. (2023); Chin et al. (2023); Tsai et al. (2023) have shown that T2I models are still vulnerable to prompt perturbations, which make these models disrupted to generate inappropriate content. Therefore, further defensive methods have been proposed, such as rejecting the generation of toxic images by detection Rando et al. (2022), guiding the model to generate safe content through embedding blocking Rombach et al. (2022) or fine-tuning Gandikota et al. (2023). Though methods to some extent alleviate the generation of harmful content, directly rejecting images can affect user usability, and optimizing the model would require obtaining the internal structure of T2I models and could be time-consuming, which lacks universality in real-world scenarios.
Recent works Wang et al. (2023); Diao et al. (2023); Hao et al. (2023) show that prompt engineering has been used to improve the performance of prompt-based models. Some works Diao et al. (2023) enhance the capability of large language models through prompt modification in text generation, math problem solving, code generation and etc. Other works Hao et al. (2023) use prompt optimization to improve T2I generation. Prompt optimization is regarded as a general method to improve prompt-based models without changing the parameters of the corresponding models. Thus, it is promising to improve the safety of T2I models through prompt engineering, i.e., revising the toxic prompt so that T2I models can generate safe images that preserve the appropriate portion of the user’s prompt (preserve the appropriate semantics of the toxic prompt). Figure 1 shows the case of a toxic prompt with the corresponding modified one by our fine-tuned model. As shown in the figure, with the revised prompt, the generated image does not contain unsafe content meanwhile is semantically close to the image generated by the toxic prompt. However, there is no work in this promising direction.
@@ 2 Related Work
Prompt engineering can be categorized into three applications for foundation models: adversarial attack Xu et al. (2022), prompt tuning, and prompt optimization. By using character-level Ebrahimi et al. (2018), word-level Garg and Ramakrishnan (2020), and sentence-level Zhao et al. (2017) perturbation on prompts, attackers can launch adversarial attacks on foundation models to mislead. Prompt tuning Jia et al. (2022) is used to transform downstream tasks into pre-training tasks through constructing templates and fine-tuning models to achieve few-shot learning. Prompt optimization aims to optimize the prompt to improve the performance of prompt-based models  Hao et al. (2023); Betker et al. (2023). For example, Promptist Hao et al. (2023) trains a language model to optimize original input prompts to generate human-preferred prompts. Prompt optimization has shown its efficiency and effectiveness in enhancing the capabilities of foundation models. In this work, we study a novel problem of prompt optimizer for T2I models to generate safe images.
@@ 3 Proposed Framework
Our framework is inspired by Promptist Hao et al. (2023). The difference is that our framework aims to produce safe prompts for T2I generation. After a user inputs a toxic prompt for the T2I generation, POSI automatically outputs the modified prompt to avoid generating inappropriate images while preserving the appropriate portion of the user’s prompt (i.e., maintaining text alignment). An illustration of the proposed framework is shown in Figure 2. Due to the absence of a publicly available toxic-clean prompts pair dataset, we first produce a set of toxic-clean prompt pairs in Section 3.1. Then we use them to conduct supervised fine-tuning (SFT) in Section 3.2 to give the model the basic ability to turn toxic prompts into clean prompts. SFT can be considered a warm-up phase, hence the effectiveness of the supervised fine-tuned model is generally moderate. To enhance the model’s performance, we further perform proximal policy optimization in Section 3.4 to maximize the reward score we design in Section 3.3. It can reduce the inappropriateness of the generated images while not significantly affecting the text alignment. Next, we give the details.
@@ 3.1 Toxic-Clean Prompt Construction
In order to give the prompt optimizer, generally implemented using as a Language Model (LM), the basic ability to modify prompts to prevent T2I models from creating inappropriate images, we need to construct a dataset containing clean-toxic prompt pairs for SFT. However, manually preparing a large number of clean-toxic prompt pairs is time-consuming. As large language models have shown great ability in following few-shot examples for text generation Brown (2020); Ouyang et al. (2022), we rely on large language models (LLMs) for generating large-scale toxic-clean pairs. We manually craft a small number of high-quality toxic-clean prompt pairs. The clean prompts are designed to effectively reduce the likelihood of generating inappropriate images while maintaining good text alignment. We then utilize these pairs as few-shot examples to ask an LLM to rewrite toxic prompts to clean prompts, thereby constructing a dataset.
Specifically, we first collect some toxic prompts from I2P dataset Schramowski et al. (2023). Then we utilize a LLM GPT-3.5 Turbo through few-shot learning to obtain the corresponding clean prompts. We denote the toxic-clean prompt pairs as $D_{SFT}=\{(x,x^{\prime})\}$, where $x$ means the original toxic prompts and $x^{\prime}$ stands for prompts modified by GPT-3.5 Turbo. The selection of GPT-3.5 Turbo is predicated on its favorable balance between performance efficacy and cost-effectiveness, relative to alternative models. The instructions we use to process toxic prompts are detailed in Appendix C.
@@ 3.2 Supervised Fine-tuning
With the toxic-clean prompts, we can now train a prompt-optimizer to let it have the basic ability of toxic prompt rewriting. Let $\pi_{\theta}$ denote the prompt optimizer that we want to train during SFT. Note that it can be any pre-trained LM. The training objective in SFT is to optimize the following loss function:
@@ 3.3 Reward Score
In order for the prompt optimizer to have the ability to rewrite toxic prompts that can generate safe and semantic-preserving images, we need to define the reward based on the modified prompt for PPO. Specifically, the modified prompts are evaluated from two views: toxicity and text alignment, where toxicity measures the probability of generated images containing inappropriate content, and text alignment measures the similarity of the generated images to the text itself.
@@
To ensure that the images generated by the modified prompts still have text alignment with the original prompts, we need to quantify the relevance between the images and the original input prompts. Specifically, similar to Promptist Hao et al. (2023), we calculate the CLIP Radford et al. (2021) scores to measure how close the generated images ${}_{x^{\prime}}$ conditioned on modified prompt $x^{\prime}$ and $x^{\prime}$ are. The alignment score can be defined as:
$S_{alt}(x^{\prime})=\mathbb{E}_{i_{x^{\prime}}\sim G(x^{\prime})}\min(0.31,f_{CLIP}(x^{\prime},i_{x^{\prime}}))$ (3)
where $f_{CLIP}(\cdot,\cdot)$ is CLIP similarity function and $i_{x^{\prime}}$ denotes the image generated by the T2I model $G$ conditioned on the the modified prompt $x^{\prime}$. The average CLIP Score of SD on I2P is around 0.3. However, we found that excessively high text alignment can impair the model’s ability to reduce the generation of inappropriate images and lead to very unstable training. Hence, we set the maximum reward for text alignment to 0.31 to ensure that while minimizing the possibility of generating inappropriate images, we maintain text alignment as close as possible to the original model. Note that we do not calculate the CLIP similarity score based on $i_{x^{\prime}}$ and original prompt $x$ directly. The reason is that our modified prompt during the early training stage remains close to the original prompt while being safer. Therefore, utilizing $x^{\prime}$ not only enhances training stability but also helps maintain text alignment with the normal part of the original toxic prompt.
We use $\pi_{\phi}$ to denote the policy model to be trained during Reinforcement Learning (RL) training and $\pi_{SFT}$ to denote the supervised fine-tuned model in Section 3.2. Following previous methods Ouyang et al. (2022); Hao et al. (2023), we also use an additional KL penalty term between $\pi_{\phi}$ and $\pi_{SFT}$ with coefficient $\beta$. This is to prevent the policy model from producing meaningless prompts in pursuit of higher rewards.
@@
$R(x,x^{\prime})=S_{toxic}+S_{alt}-\beta\mathrm{log}\frac{\pi_{\phi}(x^{\prime}|x)}{\pi_{SFT}(x^{\prime}|x)}$ (4)
@@ 3.4 Proximal Policy Optimization
With the reward score measuring both toxicity and text alignment, following Promptist Hao et al. (2023), we propose to enhance our model by employing Proximal Policy Optimization (PPO) Schulman et al. (2017) during RL training for two reasons: (i) PPO has been empirically demonstrated to be data-efficient and effective Schulman et al. (2017); Hao et al. (2023); and (ii) We could compute the reward directly from the images produced by the T2I model to conduct PPO, without requiring knowledge of the T2I model’s internal architecture. Specifically, we initialize the parameters of $\pi_{\phi}$ by $\pi_{SFT}$. We then optimize $\pi_{\phi}$ by optimizing the following objective function in RL training over the training set $D_{PPO}=\{x\}$ as:
@@ 4 Experiments
Figure 4: Examples of various methods to prevent the generation of inappropriate content on SD v1.4. Some images were manually blurred after being generated for display purposes. The prompts and their corresponding seeds and guidance scales are in Appendix A.3.
@@ 4.1 Datasets
Firstly, we extract 50 prompts from each of the 6 categories in I2P Schramowski et al. (2023), namely sexual, harassment, self-harm, illegal activity, shocking, and violence, forming an evaluation dataset. The remaining prompts in I2P are split into two parts for the SFT and PPO stages, respectively. We employ the method in section 3.1 to process toxic prompts in I2P for SFT, to create a toxic-clean prompt pairs dataset for SFT. We also use Template prompts Qu et al. (2023) as another evaluation dataset. Template prompts is a manually created prompt dataset, where phrases are filled in a fixed prompt template. Prompts in it have a high probability of causing SD to generate inappropriate images. Table 1 summarizes these datasets. Note that we originally selected 3,561 prompts from I2P for SFT, but due to some prompts being rejected by GPT, we ended up with a total of 3,390 pairs.
@@ 4.2 Baseline Methods
To test the results on SD, we choose several concept removal methods that can remove inappropriate content. For guidance-based methods, We choose SLD Schramowski et al. (2023) with four different parameter settings (SLD-Weak, SLD-Medium, SLD-Strong, and SLD-Max) and SD-NP Rombach et al. (2022). For fine-tuning-based methods, we choose ESD Gandikota et al. (2023). For ESD, we fine-tune the non-cross-attention with the strength of negative guidance set as 1 (ESD-u-1). We use the prompt “hate, harassment, violence, suffering, humiliation, harm, suicide, sexual, nudity, bodily fluids, blood, obscene gestures, illegal activity, drug use, theft, vandalism, weapons, child abuse, brutality, cruelty” Schramowski et al. (2023) for SD-NP and training ESD-u-1. Note that ESD is implemented in SD v1.4, so we only present the results of ESD on SD v1.4. The implementations of all baselines are based on their official codes.
@@ 4.3 Settings
As for our framework, for the language model, we use LLaMA Touvron et al. (2023) with 7B parameters. When calculating the alignment score, we use the CLIP based on ViT-B/32. We set the T2I model $G$ in Eq. 2 as SD v1.4. We use LoRA Hu et al. (2022) for training in both the SFT and PPO phases. We set the LoRA target as q_proj and v_proj. LoRA rank is set to 8. During SFT phase, we use a batch size of 4, a learning rate of 5e-5, and a gradient accumulation step of 4. We fine-tune the model for 3 epochs (636 steps in total). During PPO phase, we use a batch size of 4, an initial learning rate of 1.9e-5, and a gradient accumulation step of 4. We fine-tune the model for 1 epoch (52 steps in total). Similar to Promptist Hao et al. (2023), to reduce variance, we create 10 images for each prompt to get the average reward. Our code is built based on LLaMA-Factory Zheng et al. (2024).